OpenStudy (anonymous):

what is the slope intercept equation of the line ? y-intercept (0, -2) slope -1

OpenStudy (anonymous):

is it -x -2?

OpenStudy (abb0t):

\[(y-y_1)=m(x-x_1)\] rearrange in the form y = mx + b

OpenStudy (abb0t):

\[y_1 = y-value\] \[x_1 = x-value\] m = slope

OpenStudy (anonymous):

....

OpenStudy (anonymous):

b = -2, m = -1 just plug into y = mx + b